OpenStudy (sohailiftikhar):

The shape obtained by rotating y=sq root (a^2 - x^2) about x-axis is called A- circle B- semi circle C- sphere D- hemisphere E- parabola F- No definite shape

OpenStudy (irishboy123):

go down this path and see where it leads/ what you recognise: \(y=\sqrt{ a^2 - x^2} \) \(y^2=\ a^2 - x^2 \) etc....

OpenStudy (sohailiftikhar):

can you draw daigrame to elaborate this please ..?

OpenStudy (irishboy123):

i can't do it for you but we can take the equation further: \(x^2 + y^2= a^2 \) does that mean anything to you?

OpenStudy (sohailiftikhar):

yes can explain it with different method its up to you ..thank you for explaining it ..please elaborate it ..

OpenStudy (irishboy123):

what shape does that equation give in the xy plane...?

OpenStudy (sohailiftikhar):

I think semi circule...